Cos 4x = 1-8sin square x cos square x prove that the R.H.S

Grade:11

1 Answers

Arun
5 years ago
 

you can solve 

cos4x as

cos2(2x) =1-2sin2(2x)

             =1-2(2sinx . cosx)2                          { sin2x = 2sinx. cosx}

             =1-2(4sin2x.cos2x)

            =1-8sin2x.cos2x

LHS=RHS